# Low-resolution descriptions of model neural activity reveal hidden features and underlying system properties

### v2

@@ I Introduction
Here, we apply this idea to the case of a Hopfield model. While the MEOW approach only takes the time series of the neuron states as an input, applying the protocol in such a context enables us to directly relate the outcomes of the analysis to the structure and interactions of the model. Hence, this allows us to validate the workflow in a controlled case where the details of the system are known, with the perspective of employing MEOW in more complex scenarios in which only the observations of the network states are available, while the underlying generative process is not.
The MEOW analysis enables us to highlight a number of system properties related to the level of detail at which the Hopfield network is observed: in particular, we identify three distinct regimes in which qualitatively different groups of neurons are pinpointed as informative depending on the resolution level of the coarse description, the latter being roughly related to the number of constituents retained in the simplified picture. These results highlight the tight connection between the level of detail at which a system is described and the amount and quality of information that its analysis can reveal.
@@ II.1 The Hopfield model
$\mu=1,...,p,$ (1)
which are usually assumed to be independently distributed random variables, with
@@
where $\delta(\cdot,\cdot)$ represents a Kronecker delta.
@@
Despite its relatively simple mathematical formulation, the Hopfield model exhibits an extremely rich phenomenology. Indeed, it has been proven Amit_2_saturation that the efficacy of the system in retrieving the encoded memories along its dynamics is critically dependent on the amount of noise—that is, the temperature $T$ in Eq. 4—and on the ratio $\alpha=p/N$ between the number of stored patterns and the size of the network. The phase diagram of the system as a function of these parameters, sketched in Fig. 2, is characterised by a paramagnetic and a spin glass phase (above the temperature curves labelled with $T_{g}$ and $T_{M}$, respectively) in which the network is not able to retrieve any memory, as well as a retrieval phase (below $T_{M}$) where the embedded patterns appear as thermodynamic metastable (below $T_{M}$ and above $T_{C}$) or stable (below $T_{C}$) states.
@@ III.1 Matrix reconstruction and bias detection for a N=10 Hopfield model
We thus begin to examine the properties of the Hopfield model through the lenses of mapping entropy considering a small network of $N=10$ at $T=0$, with an amount of stored patterns in the range $p\in\{1,..,5\}$. As anticipated, the relatively small total number of reduced representations (irrespectively of the resolution level $n_{cg}$) that can be designed to describe a network of this size, $n_{map}=2^{10}-1=1023$, allows us to perform an exhaustive exploration of all of them for every value of $p$ investigated; the state configurations sampled by each setup in the course of its simulation were provided as input to the EXCOGITO software for the MEOW analysis. The obtained results will thus constitute a benchmark for the following analysis on a larger system discussed in Sec. III.2.
@@ III.1.1 Properties of the minima of the mapping entropy and their relation with the memory patterns
At the same time, by further analysing the behaviour of the mapping entropy as a function of the resolution at a fixed degree of detail, a closer inspection of Fig. 3 reveals that, despite the overall linearity of $S_{map}$ in $H_{S}$, for several values of $n_{cg}$ the reduced representations of the system split into distinct clusters separated by gaps in mapping entropy. This feature is particularly evident in the case of $p=2$ and $p=3$ (resp. Fig. 3b and 3c), while it is almost absent for the other amounts of stored patterns. One is then naturally led to ask why such a “classification” of decimation mappings appears; critically, the reason for this is to be found in the structure of the synaptic weight matrix $J_{ij}$ characterising the interactions among the network’s constituents. To better clarify this point, we focus on the simplest case with $p=2$ retaining $n_{cg}=3$ sites—for which a zoomed version of the results presented in Fig. 3b is reported in Fig. 4; furthermore, we first consider a Hopfield network with two “biased” patterns, namely
@@ IV Conclusions
In this work we carried out a study of the Hopfield model through the lenses of information-theoretic measures, expanding on the work done by Giulini and coworkers HGP and Roudi and coworkers On_sampling_and_modeling_Marsili ; Marsili_1 . Specifically, we investigated the properties of low-resolution representations of the model making use of the mapping entropy optimization workflow (MEOW) and an analysis based on resolution, a measure of the level of detail inherent in a dataset of coarse representations of a system’s states. Applying these tools to the configurations sampled in a number of simulations of different Hopfield networks, we observed that the low-resolution mappings with the lowest values of mapping entropy reveal information about intrinsic properties of the reference system.
@@
The results presented in this work have highlighted nontrivial relations between the level of detail of the low-resolution representation at which the system is inspected and the underlying generative process, specifically the interaction matrix of the model. The analysis has focused on particular realisation of the Hopfield network, with the aim of establishing a link the most direct possible between a case-specific system and its optimal low-resolution mappings. The picture that emerges thus provides promising evidence that such link can be leveraged to extract nontrivial information about the properties of systems when only part of them is accessible to inspection. Further work is required to assess the generality of the observed features; in particular, a broad-range campaign of simulations and analysis is in order to characterise the typical expected behaviour in various setups, e.g., averaging over the quenched disorder for different degrees of structure in the statistical properties of the stored patterns.
In conclusion, this work has shown that it is possible to gather useful knowledge about a model neural network through the analysis of the information content of its low-resolution representations; in particular, we have seen that the elements of the network that are pinpointed as the most informative ones depend on the specific decimated representation resolution level at which the system is described: this result can serve as a guide in the study of complex systems as well as in the construction of effective models of the latter, and paves the way to the development of a semi-automated protocol for the analysis of limited data gathered from systems composed by a large number of constituents.
V Acknowledgments
@@
VI Author contributions

### v3

@@ I Introduction
We here apply this idea to the case of a Hopfield model. While the MEOW approach only takes the time series of the neuron states as an input, applying the protocol in such a context enables us to directly relate the outcomes of the analysis to the structure and interactions of the model. Hence, this allows us to validate the workflow in a controlled case where the details of the system are known, focussing on the effect that specific realisations of the memory patters have on the resulting emergent behaviour of the network. This analysis paves the way to the application of MEOW in more complex scenarios in which only the observations of the network states are available, while the underlying generative process is not.
The MEOW protocol enables us to highlight a number of system properties related to the level of detail at which the Hopfield network is observed: in particular, we identify three distinct regimes in which qualitatively different groups of neurons are pinpointed as informative depending on the resolution level of the coarse description, the latter being roughly related to the number of constituents retained in the simplified picture. These results highlight the tight connection between the level of detail at which a system is described and the amount and quality of information that its analysis can reveal.
@@ II.1 The Hopfield model
$\mu=1,...,p.$ (1)
A typical statistical-mechanical analysis of the model in the thermodynamic limit relies on the assumption that the memories are independently distributed random variables, with
@@
where $\delta(\cdot,\cdot)$ represents a Kronecker delta. This, however, is not a necessary requirement, and the memory patterns can indeed entail nontrivial statistical properties.
@@
Despite its relatively simple mathematical formulation, the Hopfield model exhibits an extremely rich phenomenology. Indeed, it has been proven Amit_2_saturation that the efficacy of the system in retrieving the encoded memories along its dynamics is critically dependent on the amount of noise—that is, the temperature $T$ in Eq. 4—and on the ratio $\alpha=p/N$ between the number of stored patterns and the size of the network. In the thermodynamic limit, the phase diagram of the system as a function of these parameters, sketched in Fig. 2, is characterised by a paramagnetic and a spin glass phase (above the temperature curves labelled with $T_{g}$ and $T_{M}$, respectively) in which the network is not able to retrieve any memory, as well as a retrieval phase (below $T_{M}$) where the embedded patterns appear as thermodynamic metastable (below $T_{M}$ and above $T_{C}$) or stable (below $T_{C}$) states.
@@ III.1 Matrix reconstruction and bias detection for a N=10 Hopfield model
We thus begin to examine the properties of the Hopfield model through the lenses of mapping entropy considering a specific realisation of a small network of $N=10$ at $T=0$, with an amount of randomly-generated stored patterns in the range $p\in\{1,..,5\}$. As anticipated, the relatively small total number of reduced representations (irrespectively of the resolution level $n_{cg}$) that can be designed to describe a network of this size, $n_{map}=2^{10}-1=1023$, allows us to perform an exhaustive exploration of all of them for every value of $p$ investigated; the state configurations sampled by each setup in the course of its simulation were provided as input to the EXCOGITO software for the MEOW analysis. The obtained results will thus constitute a benchmark for the following analysis on a larger system discussed in Sec. III.2.
@@ III.1.1 Properties of the minima of the mapping entropy and their relation with the memory patterns
At the same time, by further analysing the behaviour of the mapping entropy as a function of the resolution at a fixed degree of detail, a closer inspection of Fig. 3 reveals that, despite the overall linearity of $S_{map}$ in $H_{S}$, for several values of $n_{cg}$ the reduced representations of the system split into distinct clusters separated by gaps in mapping entropy. This feature is particularly evident in the case of $p=2$ and $p=3$ (resp. Fig. 3b and 3c), while it is almost absent for the other amounts of stored patterns. One is then naturally led to ask why such a “classification” of decimation mappings appears; critically, the reason for this is to be found in the structure of the specific realisation of the synaptic weight matrix $J_{ij}$ characterising the interactions among the network’s constituents. To better clarify this point, we focus on the simplest case with $p=2$ retaining $n_{cg}=3$ sites—for which a zoomed version of the results presented in Fig. 3b is reported in Fig. 4; furthermore, we first consider a Hopfield network with two “biased” patterns, namely
@@ IV Conclusions
In this work we carried out a study of the Hopfield model through the lenses of information-theoretic measures, expanding on the work done by Giulini and coworkers HGP and Roudi and coworkers On_sampling_and_modeling_Marsili ; Marsili_1 . Specifically, we investigated the properties of low-resolution representations of specific realisations of the model making use of the mapping entropy optimization workflow (MEOW) and an analysis based on resolution, a measure of the level of detail inherent in a dataset of coarse representations of a system’s states. Applying these tools to the configurations sampled in a number of simulations of different Hopfield networks, we observed that the low-resolution mappings with the lowest values of mapping entropy reveal information about intrinsic properties of the reference system.
@@
The results presented in this work have highlighted nontrivial relations between the level of detail of the low-resolution representation at which the system is inspected and the underlying generative process, specifically the interaction matrix of the model. The analysis has focused on particular realisation of the Hopfield network, with the aim of establishing a link the most direct possible between a case-specific system and its optimal low-resolution mappings. The picture that emerges thus provides promising evidence that such link can be leveraged to extract nontrivial information about the properties of systems when only part of them is accessible to inspection.
In conclusion, this work has shown that it is possible to gather useful knowledge about a neural network through the analysis of the information content of its low-resolution representations; in particular, we have seen that the elements of the network that are pinpointed as the most informative ones depend on the specific decimated representation resolution level at which the system is described: this result can serve as a guide in the study of complex systems as well as in the construction of effective models of the latter, and paves the way to the development of a semi-automated protocol for the analysis of limited data gathered from systems composed by a large number of constituents.
Acknowledgments
@@
Data and software availability
Raw data produced and analyzed in this work are freely available on the Zenodo repository https://doi.org/10.5281/zenodo.13940980.
Author contributions